OpenStudy (anonymous):

Calculus 2 problem. Help please. y=e^x, y=e^2x, x=ln2, x=0 Find the area enclosed by the curves. Show steps please.

OpenStudy (dumbcow):

\[A = \int\limits_{0}^{\ln(2)} (e^{2x} - e^{x}) dx\]

OpenStudy (anonymous):

Right, but what is the actual area. That's where I am having trouble. thanks.

OpenStudy (anonymous):

The answer is supposed to be 1/2, but I am not getting that.

zepdrix (zepdrix):

Were you able to get through the integration step? Before plugging in the limits.

zepdrix (zepdrix):

Were you able to get this? :) \[\large \frac{1}{2}e^{2x}-e^x\]

OpenStudy (anonymous):

Yes, I got that part, but I did the algebra wrong, looking at it now. I got the answer now. I have a quick question, however. How do you know which function to use first in the integration process?

OpenStudy (anonymous):

And is there anyway to give more than one medal?

zepdrix (zepdrix):

No, just give to cow c: hehe

zepdrix (zepdrix):

You mean which one is being subtracted when you setup the integrand? Or do you mean which limit you're subtracting when you evaluate the integral (plugging in).

OpenStudy (anonymous):

I guess both. I thought I was okay on the integral, but I am questioning myself on it now.

zepdrix (zepdrix):